# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

### v1

@@ 3.2 Cancellation of factorial growth between coefficient functions and operators
Given the discussion in the previous subsection, it is clear from now that the factorial growths in the coefficient functions can not be canceled within the coefficient functions. Instead, they are expected to cancel with the operators. How to see the factorial growth in the operators? We first consider the alternating $C_{n,0}^{\rm asy2}$. It looks a operator contribution from $O_{n-1}$ but with $n$-bubbles. Let’ show it is indeed the case. For this we need
$\frac{1}{(p^{2})^{n-1+2}}\langle O_{n-1}\rangle_{g^{n+1}}=\frac{(-1)^{n+1}}{N(p^{2})^{n-1+2}}\int\frac{d^{2}k}{(2\pi)^{2}}(k^{2})^{n-2}g(gF(k^{2}))^{n}\ .$ (136)
@@
$\frac{1}{(p^{2})^{n-1+2}}\langle O_{n-1}\rangle_{g^{n+1}}$ $\rightarrow\frac{(-1)^{n+1}}{N(p^{2})^{n-1+2}}\frac{g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\gg\alpha m^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{4}}\ln^{n}\frac{k^{2}}{m^{2}}$
$\rightarrow\frac{(-1)^{n+1}}{N(p^{2})^{n-1+2}}\frac{g^{n+1}}{(4\pi)^{n}}\frac{n!}{m^{2}}\ .$ (138)
@@
the $k=0$ term is just the UV finite contribution calculated before, while for $k\geq 1$ one replaces $k$ bubble-chains with tadpole-type insertions. The $o_{k}$ for small $k$ contain factorial growth in $n$, but alternate between different $k$. The combined effect is that there is a factorial growth of the form $\frac{1}{m_{r}^{2}}\times\frac{m_{r}^{2}}{\mu^{2}}$ that exactly matches the $\frac{1}{\mu^{2}}$ in the Eq. (123). In conclusion, the alternating “UV renormalons” in $C_{n,0}$ has to be canceled by lower-dimension operators.
We then moves to the non-alternating factorial growths Eq. (122). These looks to operator contributions from $O_{n+1}$ but still with $n$-bubbles, thus there are UV divergences which should be renormalized according to our renormalization schemes. There is a contribution of order $m_{r}^{2}g^{n+1}$, while the logarithms could lead to $g^{n+1}m_{r}^{2}\times\frac{\mu^{2}}{m_{r}^{2}}n!=g^{n+1}\mu^{2}n!$. The fact that $\mu^{2}$ is in the numerator indicates that this is a UV renormalon for the operator. To show this, one can use the following trick. For this purpose one considers the point-splitting regularization
@@
Now, by subtracting out all the singular contributions at $\frac{1}{z^{2}}$, as well as finite contributions from the mass-insertions, one reaches the $O_{n}$. So one can investigate if these subtractions could introduce factorial growths. Since we are interested in the factorial growth that correlates with the coefficient functions for the identity operator with no mass insertions, one focus on $C_{{\rm I},0}$. In fact, one has here the bared one for the $n$ bubble-contribution
@@
it contain exactly a $s=1$ renormalon singularity that is none-alternating. In the large $n$ limit, the factorial growth after the IR subtraction is therefore
@@
$\langle O_{n+1}^{\rm asy1}\rangle=\frac{g^{n+1}}{(4\pi)^{n+1}}n!\mu^{2}\ .$ (143)
@@
Similarly, one can show that the factorial growths for the $m_{r}^{2}$ terms in the coefficient functions are canceled by two sources. First, the $n!$ growth from mass insertions within the bubble chains are canceled by the $O_{n}$ operator in $n-1$ bubbles. This time, the renormalon is proportional to $m_{r}^{2}$ and can be obtained in the point-splitting by subtracting the $m_{r}^{2}$ order coefficient function $C_{{\rm I},1}$ in the $\frac{1}{z^{4}}(C_{{\rm I},0}+m_{r}^{2}z^{2}C_{{\rm I},1})$. Second, the $n\times n!$ contribution from $m_{r}^{2}$ insertion on the $p-k$ line can be explained by the operator $O_{n}$ in $n-1$ bubbles but attached to the $m_{r}^{2}$ order coefficient function. The small-$k$ expansion of $-\frac{m_{r}^{2}}{(p-k)^{4}}$ introduces the factor of $-n^{2}m_{r}^{2}$. This combined with the operator’s $(n-1)!\mu^{2}\frac{g^{n}}{(4\pi)^{n}}$ exactly cancels the factorial growth in Eq. (134). So the cancellation of leading factorial growth between coefficient functions and operators are checked up to $m_{r}^{2}$.
To conclude, the cancellation pattern is: alternating “renomalons” cancels with lower-dimensional operators, while non-alternating renormalons cancels with higher-dimensional operators. The renormalon always cancels at the same numbers of $g$, and the cancellation is always at the same values of $s$. For example, the $s=2$ subleading renormalon corresponding to the second term in Eq. (122) would cancel with the sub-leading UV renormalon of the operator $O_{n+2}$ with $n$-bubbles (and diagrams that can be obtained by deleting bubbles but adding tadpole-type insertions).
@@ 3.3 Checking the OPE calculation with Mellin-Barnes
$\frac{\pi}{\sin\pi s_{1}}\left(\frac{g}{4\pi}\right)^{s_{1}}g\int\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{k^{2}}\left(\frac{k^{2}}{m^{2}}\right)^{s}\frac{1}{(p-k)^{2}+m^{2}}$
$=\left(\frac{g}{4\pi}\right)^{s_{1}}\frac{g}{4\sin\pi s_{1}}\int_{0<{\rm Re}(u)<s_{1}-s}\frac{du}{2\pi i}\frac{\Gamma(u)\Gamma(s-s_{1}+1)\Gamma(-s+s_{1}-u)^{2}}{\Gamma(1-u)\Gamma(s_{1}-s)}\left(\frac{p^{2}}{m^{2}}\right)^{-u}\ .$ (147)
@@
This is nothing but the $C^{B}_{n,0}$ term in Eq. (121) that contains factorial growth calculated from IR renormalized bubble-chains, at the natural renormalization scale $\mu^{2}=m^{2}$ that kills “tree-level” operators. As result, from the Mellin-Barnes one confirms the correctness of the fixed-pointed calculation.
Using Mellin Barnes, one can also verify the cancellation pattern between operators and coefficient functions. For example, the $t=1$ pole of ${\cal H}_{0,0}|_{s_{1}=n}$ cancels with the $t=1$ pole from $S_{0,n+1}\bigg{|}_{s_{1}=n}$. Namely, it is the $g^{n}$ contribution to the operator $O_{n+1}$ that cancels the leading factorial growth of $C_{n,0}$, consistent with the cancellation pattern suggested in the previous subsection. More generally, the $t=p$ pole of ${\cal H}_{0,k}\bigg{|}_{s_{1}=n}$ cancels the $t=p$ pole of $S_{0,k+n+p}\bigg{|}_{s_{1}=n}$, thus not only leading renormalon, sub-leading renormalons also cancel between coefficient functions and operators. Moreover, for each $l\geq 1$, the same cancellation pattern applies also to ${\cal H}_{l,k}$ and $S_{l,k}$.
Another important result that is most efficiently obtained from the Mellin-Barnes is to derive a closed formula for the leading large-$n$ asymptotics of the full coefficient functions. For this purpose, one first obtain the leading asymptotics for the contributions with $l$ mass insertions on the $p-k$ line, while with $k$ mass insertions within the bubble-chain
$\left(\frac{m^{2}}{g}\right)^{k+l}C_{n;l,k}^{\rm asy1}(\mu^{2}=m^{2})=$
$-\frac{m^{2}}{p^{2}}\frac{(-1)^{k+l}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k+l}\frac{(n!)^{2}(n-k-l)!\Gamma(k-l+n+2)}{\pi k!(l!)^{2}((n-l)!)^{2}}$
@@
Using the fact that the largest contribution is clearly supported in the region $n\gg k$, one can perform the large $n$ expansion to obtain
$\left(\frac{m^{2}}{g}\right)^{k}\frac{C_{n,l,k}^{\rm asy1}}{C_{n,l,0}^{\rm asy1}}\bigg{|}_{n\rightarrow\infty}=\frac{(-2)^{k}}{\Gamma(k+1)}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{k}+\frac{(-1)^{k}2^{k-1}}{\Gamma(k-1)}\frac{1}{n}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{k}+{\cal O}\left(\frac{l}{n^{2}}\right)\ .$ (153)
@@
$\left(\frac{m^{2}}{g}\right)^{m}C_{n;l,0}^{\rm asy1}(\mu^{2}=m^{2})=-\frac{1}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{l}\frac{m^{2}}{p^{2}}\frac{(-1)^{l}\Gamma(n+1)^{2}}{\Gamma(l+1)^{2}\Gamma(n-l+1)}\ .$ (154)
@@
One obtains after summing over $m$ and $k$, the following leading large $n$ asymptotics originated from the $s=1$ IR renormalons of the $C_{n}$
$C_{{\rm I},{\rm NL}}^{(n)}(\mu^{2}=m^{2},p^{2})\bigg{|}_{n\rightarrow\infty}$
$\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n+2}}\frac{m^{2}}{p^{2}}e^{-\frac{8\pi m^{2}}{g}}n!L_{n}\left(\frac{4\pi m^{2}}{g}\right)\bigg{(}1+{\cal O}\left(\frac{1}{n}\right)\bigg{)}\ .$ (155)
@@
Here $L_{n}(x)$ is the Laguerre polynomial with $L_{n}(0)=1$. The above asymptotics has been checked numerically given Eq. (3.3). For generic $\mu^{2}$, the natural generalization is
$C_{{\rm I},{\rm NL}}^{(n)}(\mu^{2},p^{2})\bigg{|}_{n\rightarrow\infty}$
$\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n+2}}\frac{\mu^{2}}{p^{2}}e^{-\frac{8\pi m_{r}^{2}}{g}}n!L_{n}\left(\frac{4\pi m_{r}^{2}}{g}\right)\bigg{(}1+{\cal O}\left(\frac{1}{n}\right)\bigg{)}\ .$ (156)
@@
Clearly, the factorial growth will never be canceled and the power expansion for the coefficient function as well as for the momentum space two-point function will not be convergent. The Eq. (3.3) and Eq. (3.3) represents one of the major result of this work. Leading asymptotics from the alternating UV renormalon that is proportional to $\frac{p^{2}}{\mu^{2}}$ can also be calculated in similar manners. Notice that the over-all factor $e^{-\frac{8\pi m_{r}^{2}}{g}}$ is non-perturbative in the coupling constant $g$. It is interesting to notice that if one introduce the Borel transform
@@
$B[t]=-\frac{\mu^{2}}{p^{2}}e^{-\frac{8\pi m_{r}^{2}}{g}}\sum_{n=0}^{\infty}t^{n}L_{n}\left(\frac{4\pi m_{r}^{2}}{g}\right)=\frac{\mu^{2}}{p^{2}}\frac{\exp\left(-\frac{4\pi m_{r}^{2}}{g}\frac{2-t}{1-t}\right)}{t-1}\ ,$ (157)
then there is a singularity at $t=1$ which is of infinite order. If we use the standard $(1\pm i0)$ Borel prescription, the Borel integral would be convergent for $\frac{g}{4\pi p^{2}}<1$. Moreover, in such case one also needs to use the same $(1\pm i0)$ Borel prescriptions not to define the operators, but to sum over infinitely many operator contributions.

### v2

@@ 3.2 Cancellation of factorial growth between coefficient functions and operators
Given the discussion in the previous subsection, it is clear from now that the factorial growths in the coefficient functions can not be canceled within the coefficient functions. Instead, they are expected to cancel with the operators. How to see the factorial growth in the operators? We first consider the alternating asymptotics of $C_{n,0}$ in Eq. (129). It looks like an operator contribution from $O_{n-1}$ but with $n$-bubbles. Let’ show it is indeed the case. For this we need
$\langle O_{n-1}\rangle_{g^{n+1}}=(-1)^{n+1}\int\frac{d^{2}k}{(2\pi)^{2}}(k^{2})^{n-2}g(gF(k^{2}))^{n}\ .$ (149)
@@
$\langle O_{n-1}\rangle_{g^{n+1}}$ $\rightarrow(-1)^{n+1}\frac{g^{n+1}}{(4\pi)^{n}}\int_{k^{2}\gg\alpha m^{2}}\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{4}}\ln^{n}\frac{k^{2}}{m^{2}}$
$\rightarrow\frac{(-1)^{n+1}g^{n+1}}{(4\pi)^{n+1}}\frac{n!}{m^{2}}\ .$ (151)
@@
the $k=0$ term is just the UV finite contribution calculated before, while for $k\geq 1$ one replaces $k$ bubble-chains with tadpole-type insertions. The $o_{k}$ for small $k$ contains factorial growth in $n$, but alternates between different $k$. The combined effect is that there is a factorial growth of the form $\frac{1}{m_{r}^{2}}\times\frac{m_{r}^{2}}{\mu^{2}}$ that exactly matches the $\frac{1}{\mu^{2}}$ in the Eq. (129). In conclusion, the alternating “UV renormalons” in $C_{n,0}$ have to be canceled by lower-dimension operators.
We then moves to the non-alternating factorial growths Eq. (128). They resemble operator contributions from $O_{n+1}$ but still with $n$-bubbles, thus there are UV divergences which should be renormalized according to our renormalization schemes. There is a contribution of the order $m_{r}^{2}g^{n+1}$, while the logarithms could lead to $g^{n+1}m_{r}^{2}\times\frac{\mu^{2}}{m_{r}^{2}}n!=g^{n+1}\mu^{2}n!$. The fact that $\mu^{2}$ is in the numerator indicates that this is a UV renormalon for the operator. To show this, one can use the following trick. Let’s consider the point-splitting regularization
@@
Now, by subtracting out all the singular contributions at $\frac{1}{z^{2}}$, as well as finite contributions from the mass-insertions, one reaches the $O_{n}$. So one can investigate if these subtractions could introduce factorial growths. Since we are interested in the factorial growth that correlates with the coefficient functions for the identity operator with no mass insertions, one focus on $C_{{\rm I},0}$. In fact, one has here the bared one for the $n$ bubble-contribution in it
@@
it contains exactly a $s=1$ renormalon singularity that is none-alternating. In the large $n$ limit, the factorial growth after the IR subtraction is therefore
@@
$\langle O_{n+1}\rangle\bigg{|}_{n\rightarrow\infty,+1}=\frac{g^{n+1}}{(4\pi)^{n+1}}n!\mu^{2}\ .$ (156)
@@
Similarly, one can show that the factorial growths for the $m_{r}^{2}$ terms in the coefficient functions are canceled by two sources. First, the $n!$ growth of $C_{n;0,1}$ from mass insertions within the bubble chains are canceled by the $O_{n}$ operator in $n-1$ bubbles. This time, the renormalon is proportional to $m_{r}^{2}$ and can be obtained in the point-splitting by subtracting the $m_{r}^{2}$ order coefficient function $C_{{\rm I},1}$ from the $\frac{1}{z^{4}}(C_{{\rm I},0}+m_{r}^{2}z^{2}C_{{\rm I},1})$. Second, the $n\times n!$ growth of $C_{n;1,0}$ from $m_{r}^{2}$ insertion on the $p-k$ line can be explained by the operator $O_{n}$ in $n-1$ bubbles but attached to the $m_{r}^{2}$ order coefficient function. Indeed, the small-$k$ expansion of $-\frac{m_{r}^{2}}{(p-k)^{4}}$ introduces the factor of $-n^{2}m_{r}^{2}$. This combined with the operator’s $(n-1)!\mu^{2}\frac{g^{n}}{(4\pi)^{n}}$ exactly cancels the factorial growth in Eq. (147). So the cancellation of leading factorial growth between coefficient functions and operators is checked up to $m_{r}^{2}$.
To conclude, the cancellation pattern is: alternating “renomalons” cancel with lower-dimensional operators, while non-alternating renormalons cancel with higher-dimensional operators. The renormalons always cancel at the same numbers of $g$, and the cancellation is always at the same values of $s$. For example, the $s=2$ subleading renormalon corresponding to the second term in Eq. (128) would cancel with the sub-leading UV renormalon of the operator $O_{n+2}$ with $n$-bubbles (and diagrams that can be obtained by deleting bubbles but adding tadpole-type insertions).
@@ 3.3 Checking the OPE calculation with Mellin-Barnes
$\frac{\pi}{\sin\pi s_{1}}\left(\frac{g}{4\pi}\right)^{s_{1}}g\int\frac{d^{2}k}{(2\pi)^{2}}\frac{1}{(k^{2})^{s_{1}}}\left(\frac{k^{2}}{m^{2}}\right)^{s}\frac{1}{(p-k)^{2}+m^{2}}$
$=\left(\frac{g}{4\pi m^{2}}\right)^{s_{1}}\frac{g}{4\sin\pi s_{1}}\int_{0<{\rm Re}(u)<s_{1}-s}\frac{du}{2\pi i}\frac{\Gamma(u)\Gamma(s-s_{1}+1)\Gamma(-s+s_{1}-u)^{2}}{\Gamma(1-u)\Gamma(s_{1}-s)}\left(\frac{p^{2}}{m^{2}}\right)^{-u}\ .$ (160)
@@
This is nothing but the $C^{B}_{n,0}$ term in Eq. (127) that contains factorial growth calculated from IR renormalized bubble-chains, at the natural renormalization scale $\mu^{2}=m^{2}$ that kills “tree-level” operators. As a result, from the Mellin-Barnes one confirms the correctness of the fixed-pointed calculation.
Using Mellin Barnes, one can also verify the cancellation pattern between operators and coefficient functions. In fact, this simply corresponds to the cancellation of $t\geq 1$ poles after picking up $s_{1}\in Z_{\geq 0}$ and implies that the $(1\pm i0)t$ prescriptions introduced separately to the functions $\{{\cal H}_{0,k}\}_{k\geq 0}$ and $\{S_{0,k}\}_{k\geq 1}$ cancel with each-other. For example, the $t=1$ pole of ${\cal H}_{0,0}|_{s_{1}=n}$ cancels with the $t=1$ pole from $S_{0,n+1}\bigg{|}_{s_{1}=n}$. Namely, it is the $g^{n+1}$ contribution to the operator $O_{n+1}$ that cancels the leading factorial growth of $C_{n,0}$, consistent with the cancellation pattern suggested in the previous subsection. More generally, the $t=p$ pole of ${\cal H}_{0,k}\bigg{|}_{s_{1}=n}$ cancels the $t=p$ pole of $S_{0,k+n+p}\bigg{|}_{s_{1}=n}$, thus not only leading renormalon, sub-leading renormalons also cancel between coefficient functions and operators. Moreover, for each $l\geq 1$, the same cancellation pattern applies also to ${\cal H}_{l,k}$ and $S_{l,k}$.
3.4 Leading large $n$ factorial asymptotics
Here we should mention that an important result that is most efficiently obtained from the Mellin-Barnes trick is to derive a closed formula for the leading large-$n$ asymptotics of the full coefficient functions. For this purpose, first notice that the method discussed in the previous subsection allows to obtain the following leading IR-renormalon asymptotics for the contributions with $l$ mass insertions on the $p-k$ line, while with $k$ mass insertions within the bubble-chain
$\left(\frac{m^{2}}{g}\right)^{k+l}C_{n;l,k}(\mu^{2}=m^{2})\bigg{|}_{n\rightarrow\infty,+1}=$
$-\frac{m^{2}}{p^{2}}\frac{(-1)^{k+l}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k+l}\frac{(n!)^{2}(n-k-l)!\Gamma(k-l+n+2)}{k!(l!)^{2}((n-l)!)^{2}}$
@@ 3.4 Leading large $n$ factorial asymptotics
Corrections to the above are exponentially small in $n$. Using the fact that the largest contribution is clearly supported in the region $n\gg k$, one can perform the large $n$ expansion to obtain
$\left(\frac{m^{2}}{g}\right)^{k}\frac{C_{n;l,k}}{C_{n;l,0}}\bigg{|}_{n\rightarrow\infty,+1}=\frac{(-2)^{k}}{\Gamma(k+1)}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{k}+\frac{(-1)^{k}2^{k-1}}{\Gamma(k-1)}\frac{1}{n}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{k}+{\cal O}\left(\frac{l}{n^{2}}\right)\ .$ (166)
@@
$\left(\frac{m^{2}}{g}\right)^{l}C_{n;l,0}(\mu^{2}=m^{2})\bigg{|}_{n\rightarrow\infty,+1}=-\frac{1}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g^{2}}\right)^{l}\frac{m^{2}}{p^{2}}\frac{(-1)^{l}\Gamma(n+1)^{2}}{\Gamma(l+1)^{2}\Gamma(n-l+1)}\ ,$ (167)
@@
one obtains after summing over $l$ and $k$, the following leading large $n$ asymptotics originated from the $s=1$ IR renormalons of the $C_{n}$
$C_{{\rm I},{\rm NL}}^{(n)}(\mu^{2}=m^{2},p^{2})\bigg{|}_{n\rightarrow\infty,+1}$
$\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n}}\frac{m^{2}}{p^{2}}e^{-\frac{8\pi m^{2}}{g}}n!L_{n}\left(\frac{4\pi m^{2}}{g}\right)\bigg{(}1+\left(\frac{4\pi m^{2}}{g}\right)^{2}\frac{2}{n}+{\cal O}\left(\frac{1}{n^{\frac{3}{2}}}\right)\bigg{)}\ .$ (168)
@@
Here $L_{n}$ is the Laguerre polynomial with $L_{n}(0)=1$. The above asymptotics has been checked numerically given Eq. (3.4). For generic $\mu^{2}$, the natural generalization is
$C_{{\rm I},{\rm NL}}^{(n)}(\mu^{2},p^{2})\bigg{|}_{n\rightarrow\infty,+1}$
$\rightarrow-\frac{g^{n+1}}{(4\pi)^{n+1}(p^{2})^{n}}\frac{\mu^{2}}{p^{2}}e^{-\frac{8\pi m_{r}^{2}}{g}}n!L_{n}\left(\frac{4\pi m_{r}^{2}}{g}\right)\bigg{(}1+\left(\frac{4\pi m_{r}^{2}}{g}\right)^{2}\frac{2}{n}+{\cal O}\left(\frac{1}{n^{\frac{3}{2}}}\right)\bigg{)}\ .$ (169)
@@
Clearly, the factorial growth will never be canceled and the power expansion for the coefficient function as well as for the momentum space two-point function will not be convergent. Notice that the over-all factor $e^{-\frac{8\pi m_{r}^{2}}{g}}$ is non-perturbative in the coupling constant $g$.
Similarly, at $\mu^{2}=m^{2}$, from the Mellin Barnes one can derive the following leading non-alternating factorial asymptotics up to subleading renormalons for the operator condensates
@@
$\bigg{\langle}O_{n;k}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty,+1}$ $\rightarrow\frac{(-1)^{k+1}g^{n+1}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k}\frac{(n-k)!\Gamma(k+n)}{\Gamma(k)}$
×$\times\,_{2}\tilde{F}_{1}(1-k,-k+n-1;n+1;-1)\ ,$ (170)
where $O_{n;k}$ denotes the contribution to the operator $O_{n}$ in the $n-k$ bubble diagram with $k\geq 1$. For $k\leq 0$, there is no non-alternating factorial growth, consistent with the UV renormalon interpretation for operators. Combining Eq. (3.4) and Eq. (3.4), one can see that the leading factorial growth of the operators $O_{n+1}$ in the $n-k$ bubble diagram exactly cancel that of $C_{n;0,k}$ with $k$ mass insertions within the $n-k$ bubbles. Now, if there are $l$ mass insertions on the $p-k$ line, one should expect that the $O_{n+1-l;k+1}$ combined with the $(m^{2})^{l}$ term in the coefficient function cancel with $C_{n;l,k}$. Indeed, using results in Appendix. B, one obtains
$\frac{(-1)^{l}m^{2}}{(p^{2})^{n+1}}\frac{g^{n+1}}{(4\pi)^{n+1}}\left(\frac{4\pi m^{2}}{g}\right)^{k+l}\frac{(n-l+k)!^{2}}{l!^{2}(n-l)!^{2}}\times\frac{(n-l-k)!\Gamma(n-l+k+2)}{k!}$
×$\times\,_{2}\tilde{F}_{1}(-k,n-k-l-1;n-l+2;-1)\ ,$ (171)
which exactly cancels Eq. (3.4) multiplying the overall factor $\frac{g^{n+1}}{(p^{2})^{n}}$ for $C_{n}$. This is completely consistent with the general cancellation pattern. Moreover, by summing over $k\geq 0$, the leading non-alternating factorial growth for the operator condensate reads
$\bigg{\langle}O_{n}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty,+1}\rightarrow\frac{m^{2}g^{n}}{(4\pi)^{n}}(n-1)!e^{-\frac{8\pi m^{2}}{g}}\bigg{(}1+\left(\frac{4\pi m^{2}}{g}\right)^{2}\frac{2}{n-1}+{\cal O}\left(\frac{1}{n^{\frac{3}{2}}}\right)\bigg{)}\ .$ (172)
Again, the non-alternating factorial asymptotics is exponentially small in $g$. The Eq. (3.4), Eq. (3.4) and Eq. (172) represent one of the major results of this work.
Leading asymptotics from the alternating renormalons can also be calculated in similar manners. For example, at $\mu=m$ one has
$\bigg{\langle}O_{n;-1+k}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty;-1}\rightarrow\frac{(-1)^{n}g^{n+2}}{(4\pi)^{n+2}}\left(\frac{4\pi m^{2}}{g}\right)^{k}\frac{(-k+n+1)!\Gamma(k+n-1)}{k!}$
×$\times\,_{2}\tilde{F}_{1}(-k,n-k;n-1;-1)\ .$ (173)
In particular, for $k=0$ the above exactly reduces to the Eq. (151) generated in the absolutely convergent $n+1$ bubble diagram. Notice that for $k\leq-1$ there are no leading alternating renormalon, while the $s_{1}=-q$ subleading renormalon persists up to $k=-q+1$. Summing over $k$ one obtains
$\bigg{\langle}O_{n}(\mu=m)\bigg{\rangle}_{n\rightarrow\infty;-1}$
$\rightarrow\frac{(-1)^{n}g^{n+2}}{(4\pi)^{n+2}m^{2}}(n+1)!e^{\frac{8\pi m^{2}}{g}}\bigg{(}1-\left(\frac{4\pi m^{2}}{g}\right)^{2}\frac{(\frac{g}{\pi m^{2}}-2)}{n}+{\cal O}\left(\frac{1}{g^{4}n^{2}}\right)\bigg{)}\ .$ (174)
Strikingly, the leading alternating factorial growth is exponentially large in $g$. This is not in contradiction with the fact that in the $g\rightarrow 0$ limit there are no factorial growths, since to reach the asymptotic region one needs $n\gg\frac{1}{g^{2}}$, which moves to infinity as $g\rightarrow 0$.
Given the asymptotcis of the operators and the coefficient functions, let’s consider their effects on the power expansion. Then, when projected to any given power $n$ in $(p^{2}+m^{2})^{2}S_{\rm NL}$, in the large $n$ limit one has
$(p^{2}+m^{2})^{2}S_{\rm NL}^{(n)}\bigg{|}_{n\rightarrow\infty,+1}$
$\rightarrow m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\bigg{(}\left(\frac{\hat{g}}{p^{2}}\right)^{n}(n-1)!L_{n-1}\left(\frac{m^{2}}{\hat{g}}\right)-\left(\frac{\hat{g}}{p^{2}}\right)^{n+1}n!L_{n}\left(\frac{m^{2}}{\hat{g}}\right)\bigg{)}\ ,$ (175)
where $\hat{g}=\frac{g}{4\pi}$. Clearly, the first contribution is due to the operators $O_{n-l},l\geq 0$, while the second contribution is due to the coefficient function $C_{{\rm I},{\rm NL}}^{(n)}$. Given the factorial behavior, the power expansion clearly failed to converge. If one introduces the Borel transforms
$B_{s}[t]=m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\sum_{n=0}^{\infty}t^{n}L_{n}\left(\frac{m^{2}}{\hat{g}}\right)=m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\frac{\exp\left(-\frac{m^{2}}{\hat{g}}\frac{t}{1-t}\right)}{1-t}\ ,$ (176)
$B_{h}[t]=-\frac{\hat{g}}{p^{2}}m^{2}e^{-\frac{2m^{2}}{\hat{g}}}\sum_{n=0}^{\infty}t^{n}(n+1)L_{n+1}\left(\frac{m^{2}}{\hat{g}}\right)=-\frac{\hat{g}}{p^{2}}\frac{d}{dt}B_{s}[t]\ ,$ (177)
then there is a singularity at $t=1$ which is of infinite order. The above implies that the power-expansion for the non-trivial coefficient functions, as well as the sum over non-trivial operators are all Borel non-summable. Despite this divergence, since $B_{s}[t]$ do not grow at infinity, one clearly has for $\frac{\hat{g}}{p^{2}}>0$
[…]
Namely, the full power expansion, although divergent and Borel non-summable in the strict sense, contain no “Borel ambiguity” if only $(1\pm i0)R$ paths are allowed. In fact, this is a general property for sums of the following form $a_{k+q}x^{k+q}-a_{k}x^{k}$, as far as the Borel transform of $\{a_{k}\}$ do not grow too fast at infinity.